Sunset rents an SUV at $21.95 plus $0.23 per mile. Sunrise rents the same vehicle for $24.95 plus $0.19 per mile. For what mileage is the cost the same?
Answer by checkley75(3666)

21.95+.23X=24.95+.19X
.23X-.19X=24.95-21.95
.04X=3
X=3/.04
X=75 MILES WILL MAKE THE PRICE THE SAME FOR BOTH COMPANIES.
PROOF
21.95+.23*75=24.95+.19*75
21.95+17.25=24.95+14.25
39.20=39.20